# Probabilistic Online Event Downsampling

### v1

@@ Abstract
Event cameras capture scene changes asynchronously on a per-pixel basis, enabling extremely high temporal resolution. However, this advantage comes at the cost of high bandwidth, memory, and computational demands. To address this, prior work has explored event downsampling, but most approaches rely on fixed heuristics or threshold-based strategies, limiting their adaptability. Instead, we propose a probabilistic framework, POLED, that models event importance through an event-importance probability density function (ePDF), which can be arbitrarily defined and adapted to different applications. Our approach operates in a purely online setting, estimating event importance on-the-fly from raw event streams, enabling scene-specific adaptation. Additionally, we introduce zero-shot event downsampling, where downsampled events must remain usable for models trained on the original event stream, without task-specific adaptation. We design a contour-preserving ePDF that prioritizes structurally important events and evaluate our method across four datasets and tasks—object classification, image interpolation, surface normal estimation, and object detection—demonstrating that intelligent sampling is crucial for maintaining performance under event-budget constraints.
@@ 3.1 Problem formulation
$\Phi_{\alpha}(e_{k})$ =$=\begin{cases}e_{k},&\quad\text{if condition holds}\\ \emptyset,&\quad\text{otherwise}\end{cases}$ (4)
@@ 3.2 Online Event Downsampling
To determine whether an incoming event $e_{k}$ is accepted or rejected, we explicitly adopt an online downsampling approach, meaning that only past and present events in $\mathcal{E}_{{1}\rightarrow{k}}$ are available at any given time. In contrast, an offline setting would allow access to future events, i.e., decisions regarding $e_{k}$ could be influenced by information in $\mathcal{E}_{{1}\rightarrow{l}}$, where $l>k$. A practical example of offline downsampling occurs in methods that aggregate events within a temporal window extending beyond the current event $e_{k}$
@@
where ||$\cdot$|| ⋅ | is the dimension of the set. If at any point $r>\alpha$, the event $e_{k}$ is dropped, even if it would have otherwise been accepted by the sampling method.
@@ 3.3.2 Uniform Downsampling
$\Phi_{\alpha}(e_{k})$ =$=\begin{cases}e_{k},&\quad\text{with probability }\alpha\\ \emptyset,&\quad\text{with probability }1-\alpha\end{cases}$ (7)
@@ 3.4 Probabilistic Online Event Downsampling
$\Phi_{\alpha}(e_{k})$ =$=\begin{cases}e_{k},&\quad\text{with probability }\text{P}_{\mathcal{E}}(\mathbf{x}_{k})\\ \emptyset,&\quad\text{with probability }1-\text{P}_{\mathcal{E}}(\mathbf{x}_{k})\end{cases}$ (8)
@@
We define the $\theta$ parameters to scale and shift the sigmoid function, ensuring that no event has a probability of exactly 00 (allowing all events a chance of being selected) or exactly $1$ (allowing even the most likely events to be removed occasionally).
@@ 3.4.1 Poisson-based Event PDF
To model event importance, we define the ePDF as the probability of at least one event occurring at a given spatial location $\mathbf{x}$ within the temporal window $[(n-1)T,nT]$:
@@ 4.1 Implementation details
[…]

### v2

@@ Abstract
Event cameras capture scene changes asynchronously on a per-pixel basis, enabling extremely high temporal resolution. However, this advantage comes at the cost of high bandwidth, memory, and computational demands. To address this, prior work has explored event downsampling, but most approaches rely on fixed heuristics or threshold-based strategies, limiting their adaptability. Instead, we propose a probabilistic framework, POLED, that models event importance through an event-importance probability density function (ePDF), which can be arbitrarily defined and adapted to different applications. Our approach operates in a purely online setting, estimating event importance on-the-fly from raw event streams, enabling scene-specific adaptation. Additionally, we introduce zero-shot event downsampling, where downsampled events must remain usable for models trained on the original event stream, without task-specific adaptation. We design a contour-preserving ePDF that prioritizes structurally important events and evaluate our method across four datasets and tasks—object classification, image interpolation, surface normal estimation, and object detection—demonstrating that intelligent sampling is crucial for maintaining performance under event-budget constraints. Code available here.
@@ 3.1 Problem formulation
[…]
@@ 3.2 Online Event Downsampling
To determine whether an incoming event $e_{k}$ is accepted or rejected, we explicitly adopt an online downsampling approach, meaning that only past and present events in $\mathcal{E}_{{1}\rightarrow{k}}$ are available at any given time. In contrast, an offline setting would allow access to future events, i.e., decisions regarding $e_{k}$ could be influenced by information in $\mathcal{E}_{{1}\rightarrow{l}}$, where $l>k$. A practical example of offline downsampling occurs in methods that aggregate events within a temporal window extending beyond the current event $e_{k}$ .
@@
where $|\cdot|$ is the dimension of the set. If at any point $r>\alpha$, the event $e_{k}$ is dropped, even if it would have otherwise been accepted by the sampling method.
@@ 3.3.2 Uniform Downsampling
[…]
@@ 3.4 Probabilistic Online Event Downsampling
[…]
@@
We define the $\theta$ parameters to scale and shift the sigmoid function, ensuring that no event has a probability of exactly 0 (allowing all events a chance of being selected) or exactly $1$ (allowing even the most likely events to be removed occasionally).
@@ 3.4.1 Poisson-based Event PDF
To model event importance, we define the ePDF as the probability of at least one event occurring at a given spatial location $\mathbf{x}$ within the temporal window [(n1)T,nT][(n-1)T,nT]:
@@ 4.1 Implementation details
To enforce an online processing constraint, we sequentially load datasets and process events one-by-one. All downsampling methods operate with a temporal window of $6~\text{ms}$, approximately corresponding to 160 FPS. Downsampling is performed entirely on the CPU, while all method-related experiments were conducted on a single NVIDIA Quadro RTX 8000 GPU.